OpenStudy (anonymous):

solve log (x – 21) = 2 – log x

OpenStudy (kira_yamato):

log (x - 21) + log x = 2 log [x (x - 21)] = 2 x (x - 21) = 100 x^2 - 21x - 100 = 0 Solve the equation, reject any answers for which x ≤ 21

OpenStudy (anonymous):

\[log(x-21)=2-logx\] \[log(x-21)+logx=2\] \[log{(x-21)(x)}=2\] \[10^{log{(x-21)(x)}}=10^2\] \[(x-21)(x)=100\] \[x^2-21x=100\] \[x^2-21x-100=0\] \[x^2-25x+4x-100=0\] \[x(x-25)+4(x-25)=0\] \[(x+4)(x-25)=0\] \[x=-4 , x=25\]

OpenStudy (kira_yamato):

\[x = \frac{-(-21)±\sqrt{21^2-4(1)(-100)}}{2(1)}\] \[x = \frac{21±\sqrt{841}}{2}\] \[x = \frac{21±29}{2}\] \[x = 25 OR x=-4(NA)\]